OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x+y=7 9x+(blank) y=(blank)

Directrix (directrix):

If the system is coincident, the two lines are the same line and coincide at every point. The equation of one of them must be identical to the other after the common factors of the coefficients are divided out.

Directrix (directrix):

Your task is to put those coefficients back in. 3 x+y=7 --- Multiply each term by 3 ---> 9x + 3*3y = 3*7 9 x+(blank) y=(blank)

Directrix (directrix):

9x + 3*3y = 3*7 9 x+( 9 ) y = (blank) Finish it, okay? @DudeISuckAtThis